INVESTIGATION OF LARGE-AMPLITUDE MOTIONS OF H_{5}^{+} AND THE DYNAMICS OF THE PROTON TRANSFER BETWEEN H_{3}^{+} and H_{2}

\underline{ZHOU LIN} and ANNE B. McCOY, Department of Chemistry and Biochemistry, The Ohio State University, Columbus, OH 43210.

Protonated hydrogen dimer, or H_{5}^{+}, is the intermediate of the proton-transfer reaction between H_{3}^{+} and H_{2}. The dynamics of this reaction has been postulated to play a significant role in the non-thermal H/D and ortho-/para- ratio in the interstellar clouds. As a weakly-bound, fluxional molecular ion, H_{5}^{+} has a very rich vibrational spectrum. The largeamplitude vibrational motions of H_{5}^{+} make theoretical studies interesting but challenging. This work aims at understanding how these large-amplitude motions are reflected in the dynamics of the proton transfer between H_{3}^{+} and H_{2}, or between the deuterated analogues of these two species. The shared-proton stretch mode is closely related to the proton-transfer process and is thus of particular interest. Diffusion Monte Carlo calculations of minimum energy paths^{a} are performed for the ground state and selected excited states, in order to explore how the vibrational energetics and wavefunctions evolve as H_{5}^{+} dissociates into H_{3}^{+} and H_{2}. The effects of deuteration on the structures and properties are also investigated.

aC. E. Hinkle and A. B. McCoy, J. Phys. Chem. Lett., 1, 562 (2010)